OpenStudy (anonymous):

I solved 2 cos(2x)+3=2 using inverse and then solved the same equation using the double angle formula, why are the answers different? what is the correct method?

OpenStudy (anonymous):

what is your answer when solving as usual?

OpenStudy (anonymous):

I get 60 and 120 when I use the inverse but when I use the double angle formula I get 60 120 240 and 300

OpenStudy (mertsj):

\[\cos 2x=\frac{-1}{2}\]

OpenStudy (mertsj):

\[2x= \frac{2\pi}{3}, \frac{4\pi}{3} + 2\pi n\]

OpenStudy (anonymous):

that is what i got when i did the inverse

OpenStudy (mertsj):

\[x=\frac{\pi}{3}, \frac{2\pi}{3} + 2\pi n\]

OpenStudy (mertsj):

Which double angle formula did you use?

OpenStudy (anonymous):

i used 2cos^2(x) - 1 and then distributed and got two additional answers but I wanted only answers between 0 and 360 degrees

OpenStudy (mertsj):